OpenStudy (anonymous):

The length of a garden is 3 1/2 feet greater than it's width. The length of the garden is 15 1/2 feet. What is the width of a garden

OpenStudy (anonymous):

L = 3.5+w L=15.5

OpenStudy (anonymous):

so that w = 15.5-3.5

OpenStudy (anonymous):

Good luck^^^